x dy/dx+3x+y=0


1
Expert's answer
2020-10-08T16:17:50-0400

Given equation is "x\frac{dy}{dx}+3x+y = 0"


Equation can be written as "\frac{dy}{dx} + \frac{y}{x} = - 3"


Integrating factor is "e^{\int \frac{dx}{x} } = x"


Then multiplying by the integrating factor and integrating,


"xy = \int -3xdx"

"xy = -\frac{3x^2}{2}+C"